# Ultraviolet spectroscopy reveals a hot and luminous companion to the Be star + black hole candidate MWC 656

### v1

@@
The Galactic Be star binary MWC 656 was long considered the only known Be star + black hole (BH) system, making it a critical benchmark for models of massive binary evolution and for the expected X-ray emission of Be+BH binaries. However, recent dynamical measurements cast doubt on the presence of a BH companion. We present new multi-epoch ultraviolet spectroscopy from the Hubble Space Telescope (HST), combined with high-resolution optical spectra, to reassess the nature of the companion. The far-ultraviolet spectra reveal high-ionisation features – including prominent N v and He ii lines – which are absent in the spectra of normal Be stars and are indicative of a hot, luminous companion. Spectral modelling shows that these features cannot originate from the Be star or from an accretion disc around a compact object. Instead, we find that the data are best explained by a hot (kK), compact, hydrogen-deficient star with strong wind signatures, consistent with an intermediate-mass stripped star. Our revised orbital solution and composite spectroscopic modelling yield a companion mass of , definitively ruling out a BH and disfavouring a white dwarf. MWC 656 thus joins the growing class of Be + stripped star binaries. The system’s unusual properties – including a high companion temperature and wind strength – extend the known parameter space of such binaries. The continued absence of confirmed OBe+BH binaries in the Galaxy highlights a growing tension with population synthesis models.
@@
The population of black holes (BHs) in binaries with massive stellar companions remains poorly constrained. While numerous high-mass X-ray binaries are known (Fortin et al., 2023), only a few host dynamically confirmed BH accretors, such as Cyg X-1, LMC X-1, and M33 X-7 (e.g. Ramachandran et al., 2025, 2022; Orosz et al., 2007, 2009). Even more elusive are non-accreting BHs in detached systems, with only a single such O+BH binary currently identified in the Large Magellanic Cloud (VFTS243; Shenar et al., 2022b) and two Galactic candidate systems (HD 130298 and ALS 8814; Mahy et al., 2022; An et al., 2025, but see also El-Badry et al. 2025).
This observational scarcity contrasts with predictions from population synthesis models, which suggest that BH companions to late O- and early B-type stars should be relatively common. Models predict that up to a few percent of such stars could host BHs in detached configurations (Langer et al., 2020b; Xu et al., 2025; Schürmann+2025). These OB+BH binaries are expected to represent a natural intermediate phase in the evolution of massive binaries, en route to double compact object formation, a subset of which may merge as gravitational-wave sources.
@@
MWC 656 was reported to be the first candidate Galactic Be+BH binary by Casares et al. (2014, hereafter C14). The detection was based on optical spectroscopy that revealed He ii 4686 Å emission, interpreted as arising from an accretion disc around an unseen compact object. Using radial velocity (RV) curves of both the Be star and the He ii emission line, C14 derived a modest eccentricity () and a mass ratio , implying a companion mass of 3.8–6.9 M⊙ for an assumed Be star mass of 10–16 M⊙. The absence of spectroscopic features from a luminous companion supported the interpretation of a BH. Additional evidence came from the faint X-ray luminosity of the system (; Ribó et al., 2017) and its location in the radio/X-ray luminosity plane, which is consistent with the properties of quiescent BH binaries (Dzib et al., 2015; Ribó et al., 2017). The system was also suggested as a candidate -ray binary (Lucarelli et al., 2010), although subsequent analysis showed that the detected emission likely originates from a background quasar (Alexander and McSwain, 2015). The Be star itself has been classified as a B1.5–2 III star (C14), with an effective temperature of kK and a mass of roughly 10 M⊙ as derived by Williams et al. (2010). Its rotational velocity was measured at km s-1 (Zamanov et al., 2021).
@@
However, the interpretation of MWC 656 as a Be+BH binary has been increasingly questioned. Rivinius et al. (2024, hereafter R24) reanalysed the system using high-resolution optical spectroscopy and found lower RV amplitudes for the Be star (–15 km s-1 based on He i 6678 absorption), compared to the km s-1 derived by C14 using Fe ii emission. This discrepancy led to a revised, much lower companion mass, inconsistent with a BH and instead suggesting a hot subdwarf companion. Independent observations by Janssens et al. (2023, hereafter J23) supported this view. They inferred a mass ratio and favoured a hot subdwarf, white dwarf, or neutron star companion. The different RV amplitudes reported in these studies likely arise from the choice of spectral tracers. While He i absorption lines trace the photosphere of the Be star, Fe ii emission originates in the circumstellar disc and can vary with tidal distortion or disc asymmetries (Panoglou et al., 2016), possibly biassing velocity measurements (see also Sect. 3.2).
@@
The uncertain nature of the companion in MWC 656 motivates further investigation. Ultraviolet (UV) spectroscopy offers a powerful means to resolve this ambiguity: stripped helium stars emit most of their radiation in the UV, resulting in a more favourable flux ratio compared to optical wavelengths. This facilitates their detection in composite spectra, as demonstrated by the discovery of ten Galactic Be+subdwarf binaries using UV spectroscopy by Wang et al. (2021), and offers the potential to constrain the wind properties and ionising flux of the companion. UV observations also yield improved constraints on the properties of the Be star, in particular its mass, luminosity, and wind parameters, since optical spectra can be contaminated by emission from the circumstellar disc. A detailed characterisation of the Be star is still lacking, and disentangling its spectral contribution is essential to correctly infer the companion mass.
@@
In this paper, we combine new high-resolution optical spectroscopy with UV spectra obtained using the Hubble Space Telescope (HST). The optical data enable a refined orbital solution and updated constraints on the Be star’s parameters. The HST far-UV spectroscopy allows us to search directly for the signature of a hot, luminous companion. This paper is organised as follows: In Section 2, we describe the observations and data reduction. Section 3 presents the revised orbital solution, Section 4 details the analysis of the Be star, and Section 5 examines the UV spectra for evidence of a companion. In Section 6 we jointly model UV and optical spectra with a composite binary model. In Section 7, we discuss the nature of the companion. We place MWC 656 in the context of known post-interaction binaries in Section 8 before summarising our conclusions in Section 9.
@@ 2.1 Optical spectroscopy
We obtained nine new optical spectra of MWC 656 using the Tillinghast Reflector Echelle Spectrograph (TRES; Fűrész 2008) on the 1.5 m Tillinghast Reflector telescope at the Fred Lawrence Whipple Observatory, Mount Hopkins, Arizona. TRES has a resolving power of and covers a wavelength range from 3900 to 9100 Å. The spectra were reduced following the procedure described in Buchhave et al. (2010), including bias subtraction, flat-field correction, and wavelength calibration.
@@
We also retrieved archival high-resolution spectra from multiple facilities. These include three epochs of observations with ESPaDOnS (; Donati 2003) at the Canada-France-Hawaii Telescope (CFHT), covering –10000 Å. Each ESPaDOnS spectrum consists of eight co-added exposures taken on a single night to increase the signal-to-noise ratio (S/N). Furthermore, we retrieved 12 archival spectra obtained with the ARC Echelle Spectrograph (ARCES; Wang et al. 2003) at Apache Point Observatory. ARCES offers a resolving power of , with coverage from 3500 to 10500 Å.
We initially considered the 34 archival spectra taken with the Fibre-fed RObotic Dual-beam Optical Spectrograph (FRODOspec; Morales-Rueda et al. 2004), which were also used by C14, J23 and R24. However, as noted by J23, these data are affected by a non-linear wavelength calibration issue. Our own inspection of interstellar Ca ii H and K absorption features confirms wavelength offsets of the order of , which decrease toward longer wavelengths. J23 mitigated this by using selected wavelength regions and applying local corrections. Since the affected spectra are not essential to our analysis and residual distortions could not be fully corrected by a simple velocity shift, we excluded them to avoid introducing systematic uncertainties in our RV measurements.
All 29 retained spectra were individually continuum-normalised using cubic spline fits to line-free regions. The spline anchor points were manually selected in wavelength intervals free of strong absorption or emission. We applied barycentric corrections to all spectra and empirically estimated the S/N from the pixel-to-pixel variation in continuum regions. A summary of all optical observations, including dates and typical S/N values, is provided in Table 4 in the Appendix.
@@ 2.2 UV spectroscopy
We obtained multiepoch UV spectroscopy of MWC 656 with the HST, combining observations from two programmes (programme 17074 PI: V. Ramachandran, programme 17202 PI: K. El-Badry). The data set includes:
A far-UV spectrum obtained with the Cosmic Origins Spectrograph (COS) using segment B of the FUV channel with the G130M grating (central wavelength 1096 Å), covering 940–1090 Å and with a resolving power –.
Four far-UV spectra obtained with the Space Telescope Imaging Spectrograph (STIS) using the FUV-MAMA detector and the E140M echelle grating, providing a resolving power of over 1150–1700 Å.
@@
We retrieved fully reduced, flux- and wavelength-calibrated spectra from the MAST archive. COS data were processed using the standard CalCOS pipeline (Soderblom, 2022), while STIS data were reduced with CalSTIS (Sohn, 2019). For the STIS echelle spectra, we merged the individual orders onto a one-dimensional wavelength grid uniformly spaced in logarithmic wavelength.
@@ 3.1 Radial velocities
To measure the RVs of the Be star, we cross-correlated the observed spectra with a synthetic template spectrum. We used a model atmosphere generated with the Potsdam Wolf-Rayet (PoWR) code for OB-type stars (Gräfener et al., 2002; Hamann and Gräfener, 2003; Sander et al., 2015). This code iteratively solves the radiative transfer equation for a spherically expanding atmosphere and statistical equilibrium equations. PoWR is a state-of-the-art code in the field of analysing hot stars with winds (see, e.g., Sander et al., 2024, for a recent comparison) and has been successfully used for the spectral analysis of Be binaries with a stripped star companion (e.g. Ramachandran et al., 2023, 2024). For the cross-correlation template, we adopted a model with , , and solar metallicity from the OB-type Milky Way grid (Hainich et al., 2019).
@@
For the companion star, the He ii 4686 emission line is the only clear and reliable tracer of orbital motion in the optical due to the absence of detectable absorption features. We determined its RVs through iterative cross-correlation centred on the He ii emission region. To this end, we used the spectrum with the highest S/N as the initial template to derive relative RVs. In subsequent iterations, a new cross-correlation template was constructed by co-adding the spectra after shifting them to the rest frame according to the derived velocities. This method is based on relative shifts and does not constrain the companion’s systemic velocity, which we fitted instead during orbital analysis.
The RVs determined for both components are listed in Table 4. RVs were measured by fitting a parabola to the peak of the cross-correlation function. The uncertainty in each RV was estimated following the prescription from Zucker (2003). This approach yields an analytical estimate of the formal uncertainty based on the local shape of the CCF peak. The quoted uncertainties represent lower bounds, as they do not account for systematic effects such as template mismatch.
@@ 3.2 Orbital analysis
We inferred the orbital parameters by simultaneously fitting the RV curves of both the Be star and its companion using a nested sampling framework. Specifically, we employed the MLFriends algorithm (Buchner, 2016, 2019), implemented in the UltraNest package (Buchner, 2021), to sample the posterior probability distributions.
The orbit of a double-lined spectroscopic binary can be described by seven parameters: the RV semi-amplitudes of the two components (, ), the orbital period (), eccentricity (), argument of periastron (), systemic velocity (), and a phase parameter (), which we defined relative to a fixed reference time () as This parameterisation follows the implementation used in Blunt et al. (2020). Since the RVs of the companion were determined relative to a template and lack an absolute zero point, we introduced an additional free parameter (), which represents a constant velocity offset between the measured relative companion RVs and the barycentric velocity of the system. To account for potentially underestimated uncertainties, we also fit RV jitter terms () that are added in quadrature to the RV uncertainties determined from the CCF peak. We adopted uniform priors for all model parameters. The prior ranges were as follows: d; km s-1; ; ; km s-1; and km s-1.
@@
Our dynamical mass ratio is lower than the reported by J23, who measured km s-1 and km s-1. The discrepancy may arise from the smaller and lower-S/N data set used in that study (18 HERMES spectra compared to 29 spectra used here). C14 reported a similar companion semi-amplitude ( km s-1) but a substantially higher Be star amplitude ( km s-1), possibly due to the inclusion of FRODOspec data, which we find suffers from unreliable wavelength calibration, or because Fe ii emission lines were used as RV tracers. As noted by R24, the double-peaked emission lines from the Be star’s disc vary in peak separation, likely due to tidal distortion of the disc (Panoglou et al., 2016; Rubio et al., 2025), making them unreliable for measuring RV amplitudes.
@@ 4 Stellar parameters of the Be star
We first analysed the Be star under the assumption that it dominates the observed light, that is, a “single-star” approach treating the companion as dark. This provides a first-order characterisation of the Be star’s properties. If the companion contributes significantly, the derived luminosity and radius would be slight overestimates (though Sect. 6 shows such corrections are small). This single-star analysis establishes the foundation for identifying discrepant spectral features in Sect. 5 and for composite modelling in Sect. 6.
@@
We estimated the fundamental parameters of the Be star by simultaneously fitting diagnostic lines in the optical and UV spectra with synthetic models. We used the highest S/N normalised optical spectrum shifted to the Be star rest frame and the UV spectrum, and compared them to a grid of synthetic models from the non-LTE PoWR model atmosphere code (Gräfener et al., 2002; Hamann and Gräfener, 2003; Sander et al., 2015), convolved with instrumental broadening kernels. As a starting point, we selected models from the OB-type Milky Way grid (Hainich et al., 2019) with solar metallicity and then computed additional models varying and to refine the fit. The effective temperature of the Be star was predominantly constrained using the Si iii /Si ii and He i /Mg ii line ratios, while surface gravity was estimated from the pressure-broadened wings of the Balmer lines. We primarily used the H, H, and H lines, which are less affected by emission from the circumstellar disc, and also considered the outer wings of H, explicitly excluding the line cores in all cases.
@@
To estimate projected rotational velocity () and macroturbulent velocity (), we applied the Fourier transform and the goodness-of-fit technique implemented in the tool iacob-broad (Simón-Díaz and Herrero, 2014). We used He i absorption lines in the blue part of the optical spectra (e.g. He i 4026, 4144, 4388). We determined km s-1 and km s-1 by averaging the results over individual lines, with the standard deviation representing the uncertainties (we note that the actual macroturbulent broadening could be lower, given the degeneracy with ). These values were then used to convolve the synthetic spectra for a direct comparison with the observations.
@@
We modelled the spectral energy distribution (SED) using the flux-calibrated UV spectra and archival photometry in the U, B, V, and Gaia G bands (Jaschek and Egret, 1982; Fabricius et al., 2002; Gaia Collaboration et al., 2023). Near-infrared bands (2MASS , , and ; Cutri et al., 2003) were not included in the fit to avoid contamination by thermal emission from the circumstellar disc of the Be star, causing an infrared excess (Williams et al., 2010). We adopted the extinction law as prescribed by Fitzpatrick (1999). We simultaneously optimised the colour excess (), the total-to-selective extinction ratio (), and a luminosity () scaling factor. The model flux was scaled using a fixed distance modulus of 11.47, which corresponds to 1.97 kpc and is based on the Gaia DR3 parallax measurement ( mas; Gaia Collaboration et al. 2023; Lindegren 2020)222We corrected the parallax measurement for the Gaia parallax zero point of mas (Lindegren, 2020) but note that the resulting changes in distance modulus by 0.1 mag and system luminosity by are negligible compared to other uncertainties in the stellar parameters.. The optimal fit parameters were determined through iterative minimisation, ensuring that the model SED accurately matched both the observed spectrum and multi-band photometry.
@@
The best-fit model yields kK, , and , corresponding to a radius of . The quoted uncertainties reflect the quality of the fit and the resolution of the model grid. While the temperature is considered reliable and consistent with the spectral type B1.5–2 III, the derived surface gravity may be underestimated. Rapid rotation significantly changes the Be star’s spherical structure, causing it to flatten at the poles and bulge at the equator. This oblateness means that the stellar radius and surface gravity are no longer uniform across the stellar surface, with both differing from pole to equator. When considering the effect of the Be star’s rapid rotation ( km s-1), the dynamically corrected surface gravity (; Repolust et al., 2004) is . This ‘true ’ provides a more accurate representation of the effective surface gravity. Specifically, it accounts for the centrifugal force resulting from the stellar rotation, which directly impacts the effective gravity and is crucial to accurately deriving the stellar mass. Despite this correction, we caution that uncertainties may still be underestimated. The wind parameters of the Be star were constrained based on diagnostic lines in the UV, in particular the Si iv doublet. The absence of P Cygni-like line profiles indicates moderate wind strength, and we found a best-fit mass loss rate at fixed terminal wind velocity of km s-1. We note that non-spherical wind structures, expected in Be stars and not captured by one-dimensional models, may affect the detailed line shape and velocity centroid (e.g. Puls et al., 2008).
@@
Building on the analysis of the fundamental parameters, we further estimated the CNO surface abundances in the Be star. We found that the observed optical and UV absorption lines for these elements did not match models assuming solar abundances. Specifically, the lines for carbon and oxygen appeared weaker, while those for nitrogen appeared stronger. To better fit the observed absorption lines, we adjusted the CNO abundances in our models (see Table 3 in Appendix A). The observed spectra are better reproduced by a model with a factor of three higher nitrogen abundance and factors of three and two lower abundances for carbon and oxygen, respectively. The CNO-processed material on the surface of the Be star could be a result of its rapid rotation, or past mass transfer from the companion star (e.g. Maeder and Meynet, 2000; Langer et al., 2008, though see also Hunter et al. 2008).

### v2

@@
The Galactic Be star binary MWC 656 was long considered the only known Be star+black hole (BH) system, making it a critical benchmark for models of massive binary evolution and for the expected X-ray emission of Be+BH binaries. However, recent dynamical measurements cast doubt on the presence of a BH companion. We present new multi-epoch ultraviolet spectroscopy from the Hubble Space Telescope, combined with high-resolution optical spectra, to reassess the nature of the companion. The far-ultraviolet spectra reveal high-ionisation features – including prominent N v and He ii lines – which are absent in the spectra of normal Be stars and are indicative of a hot, luminous companion. Spectral modelling shows that these features cannot originate from the Be star or from an accretion disc around a compact object. Instead, we find that the data are best explained by a hot (kK), compact, hydrogen-deficient star with strong wind signatures, consistent with an intermediate-mass stripped star. Our revised orbital solution and composite spectroscopic modelling yield a companion mass of , definitively ruling out a BH and disfavouring a white dwarf. MWC 656 thus joins the growing class of Be + stripped star binaries. The system’s unusual properties – including a high companion temperature and wind strength – extend the known parameter space of such binaries. The continued absence of confirmed OBe+BH binaries in the Galaxy highlights a growing tension with population synthesis models.
@@
The population of black holes (BHs) in binaries with massive stellar companions remains poorly constrained. While numerous high-mass X-ray binaries are known (Fortin et al., 2023), only a few host dynamically confirmed BH accretors, such as Cyg X-1, LMC X-1, and M33 X-7 (e.g. Ramachandran et al., 2025, 2022; Orosz et al., 2007, 2009). Even more elusive are non-accreting BHs in detached systems, with only a single such O+BH binary currently identified in the Large Magellanic Cloud (VFTS243; Shenar et al., 2022b) and two Galactic candidate systems (HD 130298 and ALS 8814; Mahy et al., 2022, An et al. in prep., but see also El-Badry et al. 2025).
This observational scarcity contrasts with predictions from population synthesis models, which suggest that BH companions to late O- and early B-type stars should be relatively common. Models predict that up to a few percent of such stars could host BHs in detached configurations (Langer et al., 2020b; Xu et al., 2025; Schürmann+2025). These OB+BH binaries are expected to represent a natural intermediate phase in the evolution of massive binaries en route to double compact object formation, a subset of which may merge as gravitational-wave sources.
@@
MWC 656 was the first candidate Galactic Be+BH binary and was reported by Casares et al. (2014, hereafter C14). The detection was based on optical spectroscopy that revealed He ii 4686 Å emission, which was interpreted as arising from an accretion disc around an unseen compact object. Using the radial velocity (RV) curves of both the Be star and the He ii emission line, C14 derived a modest eccentricity () and a mass ratio () of , implying a companion mass of 3.8–6.9 M⊙ for an assumed Be star mass of 10–16 M⊙. The absence of spectroscopic features from a luminous companion supported the interpretation of a BH. Additional evidence came from the faint X-ray luminosity of the system (; Ribó et al., 2017) and its location in the radio/X-ray luminosity plane, which is consistent with the properties of quiescent BH binaries (Dzib et al., 2015; Ribó et al., 2017). The system was also suggested as a candidate -ray binary (Lucarelli et al., 2010), although subsequent analysis showed that the detected emission likely originates from a background quasar (Alexander and McSwain, 2015). The Be star itself has been classified as a B1.5–2 III star (C14), with an effective temperature () of kK and a mass of roughly 10 M⊙ as derived by Williams et al. (2010). Its rotational velocity was measured at km s-1 (Zamanov et al., 2021).
@@
However, the interpretation of MWC 656 as a Be+BH binary has been increasingly questioned. Rivinius et al. (2024, hereafter R24) re-analysed the system using high-resolution optical spectroscopy and found lower RV amplitudes for the Be star (–15 km s-1 based on He i 6678 absorption), compared to the km s-1 derived by C14 using Fe ii emission. This discrepancy led to a revised, much lower companion mass that is inconsistent with a BH and instead suggests a hot subdwarf companion. Independent observations by Janssens et al. (2023, hereafter J23) supported this view. They inferred a mass ratio and favoured a hot subdwarf, white dwarf, or neutron star companion. The different RV amplitudes reported in these studies likely arise from the choice of spectral tracers. While He i absorption lines trace the photosphere of the Be star, Fe ii emission originates in the circumstellar disc and can vary with tidal distortion or disc asymmetries (Panoglou et al., 2016), possibly biassing velocity measurements (see also Sect. 3.2).
@@
The uncertain nature of the companion in MWC 656 motivates further investigation. Ultraviolet (UV) spectroscopy offers a powerful means to resolve this ambiguity: stripped helium stars emit most of their radiation in the UV, resulting in a more favourable flux ratio compared to optical wavelengths. This facilitates their detection in composite spectra, as demonstrated by the discovery of ten Galactic Be+subdwarf binaries by Wang et al. (2021) using UV spectroscopy, and offers the potential to constrain the wind properties and ionising flux of the companion. UV observations also yield improved constraints on the properties of the Be star, in particular its mass, luminosity, and wind parameters, since optical spectra can be contaminated by emission from the circumstellar disc. A detailed characterisation of the Be star is still lacking, and disentangling its spectral contribution is essential to correctly inferring the companion mass.
@@
In this study we combined new high-resolution optical spectroscopy with UV spectra obtained using the Hubble Space Telescope (HST). The optical data enabled a refined orbital solution and updated constraints on the Be star’s parameters. The HST far-ultraviolet (FUV) spectroscopy allowed us to search directly for the signature of a hot, luminous companion. This paper is organised as follows: In Sect. 2 we describe the observations and data reduction. Section 3 presents the revised orbital solution, Sect. 4 details the analysis of the Be star, and Sect. 5 examines the UV spectra for evidence of a companion. In Sect. 6 we jointly model UV and optical spectra with a composite binary model. In Sect. 7 we discuss the nature of the companion. We place MWC 656 in the context of known post-interaction binaries in Sect. 8 before summarising our conclusions in Sect. 9.
@@ 2.1 Optical spectroscopy
We obtained nine new optical spectra of MWC 656 using the Tillinghast Reflector Echelle Spectrograph (TRES; Fűrész 2008) on the 1.5 m Tillinghast Reflector telescope at the Fred Lawrence Whipple Observatory, Mount Hopkins, Arizona. TRES has a resolving power () of and covers a wavelength range from 3900 to 9100 Å. The spectra were reduced following the procedure described in Buchhave et al. (2010), including bias subtraction, flat-field correction, and wavelength calibration.
@@
We also retrieved archival high-resolution spectra from multiple facilities. These include three epochs of observations with the Echelle SpectroPolarimetric Device for the Observation of Stars (ESPaDOnS, ; Donati 2003) at the Canada-France-Hawaii Telescope (CFHT), covering –10000 Å. Each ESPaDOnS spectrum consists of eight co-added exposures taken on a single night to increase the signal-to-noise ratio (S/N). Furthermore, we retrieved 12 archival spectra obtained with the ARC Echelle Spectrograph (ARCES; Wang et al. 2003) at Apache Point Observatory. ARCES offers a resolving power of , with coverage from 3500 to 10500 Å.
We initially considered the 34 archival spectra taken with the Fibre-fed RObotic Dual-beam Optical Spectrograph (FRODOspec; Morales-Rueda et al. 2004), which were also used by C14, J23 and R24. However, as noted by J23, these data are affected by a non-linear wavelength calibration issue. Our own inspection of interstellar Ca ii H and K absorption features confirms wavelength offsets of the order of , which decrease towards longer wavelengths. J23 mitigated this by using selected wavelength regions and applying local corrections. Since the affected spectra are not essential to our analysis and residual distortions could not be fully corrected by a simple velocity shift, we excluded them to avoid introducing systematic uncertainties in our RV measurements.
All 29 retained spectra were individually continuum-normalised using cubic spline fits to line-free regions. The spline anchor points were manually selected in wavelength intervals free of strong absorption or emission. We applied barycentric corrections to all spectra and empirically estimated the S/N value from the pixel-to-pixel variation in continuum regions. A summary of all optical observations, including dates and typical S/N values, is provided in Table 4.
@@ 2.2 UV spectroscopy
We obtained multi-epoch UV spectroscopy of MWC 656 with the HST, combining observations from two programmes (programme 17074 PI: V. Ramachandran, programme 17202 PI: K. El-Badry). The dataset includes:
A FUV spectrum obtained with the Cosmic Origins Spectrograph (COS) using segment B of the FUV channel with the G130M grating (central wavelength 1096 Å), covering 940–1090 Å and with a resolving power () of –.
Four FUV spectra obtained with the Space Telescope Imaging Spectrograph (STIS) using the FUV-MAMA detector and the E140M echelle grating, providing a resolving power of over 1150–1700 Å.
@@
We retrieved fully reduced, flux- and wavelength-calibrated spectra from the MAST archive. COS data were processed using the standard CalCOS pipeline (Soderblom, 2022), while STIS data were reduced with CalSTIS (Sohn, 2019). For the STIS echelle spectra, we merged the individual orders onto a 1D wavelength grid uniformly spaced in logarithmic wavelength.
@@ 3.1 Radial velocities
To measure the RVs of the Be star, we cross-correlated the observed spectra with a synthetic template spectrum. We used a model atmosphere generated with the Potsdam Wolf-Rayet (PoWR) code for OB-type stars (Gräfener et al., 2002; Hamann and Gräfener, 2003; Sander et al., 2015). This code iteratively solves the radiative transfer equation for a spherically expanding atmosphere and statistical equilibrium equations. PoWR is a state-of-the-art code in the field of analysing hot stars with winds (see e.g. Sander et al. 2024 for a recent comparison) and has been used successfully for the spectral analysis of Be binaries with a stripped star companion (e.g. Ramachandran et al., 2023, 2024). For the cross-correlation template, we adopted a model with , , and solar metallicity from the OB-type Milky Way grid (Hainich et al., 2019).
@@
For the companion star, the He ii 4686 emission line is the only clear and reliable tracer of orbital motion in the optical spectrum because of the absence of detectable absorption features. We determined its RVs through iterative cross-correlation centred on the He ii emission region. To this end, we used the spectrum with the highest S/N as the initial template to derive relative RVs. In subsequent iterations, a new cross-correlation template was constructed by co-adding the spectra after shifting them to the rest frame according to the derived velocities. This method is based on relative shifts and does not constrain the companion’s systemic velocity, which we fitted instead during orbital analysis.
The RVs determined for both components are listed in Table 4. RVs were measured by fitting a parabola to the peak of the cross-correlation function. The uncertainty in each RV was estimated following the prescription from Zucker (2003). This approach yields an analytical estimate of the formal uncertainty based on the local shape of the cross-correlation function peak. The quoted uncertainties represent lower bounds because they do not account for systematic effects such as template mismatch.
@@ 3.2 Orbital analysis
We inferred the orbital parameters by simultaneously fitting the RV curves of both the Be star and its companion using a nested sampling framework. Specifically, we used the MLFriends algorithm (Buchner, 2016, 2019), implemented in the package UltraNest (Buchner, 2021), to sample posterior probability distributions.
The orbit of a double-lined spectroscopic binary can be described by seven parameters: the RV semi-amplitudes of the two components (, ), the orbital period (), eccentricity (), argument of periastron (), systemic velocity (), and a phase parameter (), which we defined relative to a fixed reference time () as This parameterisation follows the implementation used in Blunt et al. (2020). Since the RVs of the companion were determined relative to a template and lack an absolute zero point, we introduced an additional free parameter (), which represents a constant velocity offset between the measured relative companion RVs and the barycentric velocity of the system. To account for potentially underestimated uncertainties, we also fitted RV jitter terms () that are added in quadrature to the RV uncertainties determined from the cross-correlation function peak. We adopted uniform priors for all model parameters. The prior ranges were as follows: d; km s-1; ; ; km s-1; and km s-1.
@@
Our dynamical mass ratio is lower than the reported by J23, who measured km s-1 and km s-1. The discrepancy may arise from the smaller and lower-S/N dataset used in that study (18 HERMES spectra compared to 29 spectra used here). C14 reported a similar companion semi-amplitude ( km s-1) but a substantially higher Be star amplitude ( km s-1), possibly due to the inclusion of FRODOspec data, which we find suffer from unreliable wavelength calibration, or because Fe ii emission lines were used as RV tracers. As noted by R24, the double-peaked emission lines from the Be star’s disc vary in peak separation, likely due to tidal distortion of the disc (Panoglou et al., 2016; Rubio et al., 2025), making them unreliable for measuring RV amplitudes.
@@ 4 Stellar parameters of the Be star
We first analysed the Be star under the assumption that it dominates the observed light, that is, a ‘single-star’ approach treating the companion as dark. This provides a first-order characterisation of the Be star’s properties. If the companion contributes significantly, the derived luminosity and radius would be slight overestimates (though Sect. 6 shows such corrections are small). This single-star analysis establishes the foundation for identifying discrepant spectral features in Sect. 5 and for composite modelling in Sect. 6.
@@
We estimated the fundamental parameters of the Be star by simultaneously fitting diagnostic lines in the optical and UV spectra with synthetic models. We used the highest S/N normalised optical spectrum shifted to the Be star rest frame and the UV spectrum, and compared them to a grid of synthetic models from the non-local thermodynamic equilibrium PoWR model atmosphere code (Gräfener et al., 2002; Hamann and Gräfener, 2003; Sander et al., 2015), convolved with instrumental broadening kernels. As a starting point, we selected models from the OB-type Milky Way grid (Hainich et al., 2019) with solar metallicity and then computed additional models varying and to refine the fit. The effective temperature of the Be star was predominantly constrained using the Si iii /Si ii and He i /Mg ii line ratios, while surface gravity was estimated from the pressure-broadened wings of the Balmer lines. We primarily used the H, H, and H lines, which are less affected by emission from the circumstellar disc, and also considered the outer wings of H, explicitly excluding the line cores in all cases.
@@
To estimate projected rotational velocity () and macroturbulent velocity (), we applied the Fourier transform and the goodness-of-fit technique implemented in the iacob-broad tool (Simón-Díaz and Herrero, 2014). We used He i absorption lines in the blue part of the optical spectra (e.g. He i 4026, 4144, 4388). We determined km s-1 and km s-1 by averaging the results over individual lines, with the standard deviation representing the uncertainties (we note that the actual macroturbulent broadening could be lower, given the degeneracy with ). These values were then used to convolve the synthetic spectra for a direct comparison with the observations.
@@
We modelled the spectral energy distribution (SED) using the flux-calibrated UV spectra and archival photometry in the U, B, V, and Gaia G bands (Jaschek and Egret, 1982; Fabricius et al., 2002; Gaia Collaboration et al., 2023). Near-infrared bands (2MASS , , and ; Cutri et al., 2003) were not included in the fit to avoid contamination by thermal emission from the circumstellar disc of the Be star, causing an infrared excess (Williams et al., 2010). We adopted the extinction law as prescribed by Fitzpatrick (1999). We simultaneously optimised the colour excess (), the total-to-selective extinction ratio (), and a luminosity () scaling factor. The model flux was scaled using a fixed distance modulus of 11.47, which corresponds to 1.97 kpc and is based on the Gaia Data Release 3 parallax measurement ( mas; Gaia Collaboration et al. 2023; Lindegren 2020)222We corrected the parallax measurement for the Gaia parallax zero point of mas (Lindegren, 2020) but note that the resulting changes in distance modulus by 0.1 mag and system luminosity by are negligible compared to other uncertainties in the stellar parameters.. The optimal fit parameters were determined through iterative minimisation, ensuring that the model SED accurately matched both the observed spectrum and multi-band photometry.
@@
The best-fit model yields kK, , and , corresponding to a radius of . The uncertainties quoted reflect the quality of the fit and the resolution of the model grid. While the temperature is considered reliable and consistent with the spectral type B1.5–2 III, the derived surface gravity may be underestimated. Rapid rotation significantly changes the Be star’s spherical structure, causing it to flatten at the poles and bulge at the equator. This oblateness means that the stellar radius and surface gravity are no longer uniform across the stellar surface, with both differing from pole to equator. When considering the effect of the Be star’s rapid rotation ( km s-1), the dynamically corrected surface gravity (; Repolust et al., 2004) is . This ‘true ’ provides a more accurate representation of the effective surface gravity. Specifically, it accounts for the centrifugal force resulting from the stellar rotation, which directly impacts the effective gravity and is crucial to accurately deriving the stellar mass. Despite this correction, we caution that uncertainties may still be underestimated. The wind parameters of the Be star were constrained on the basis of diagnostic lines in the UV, in particular the Si iv doublet. The absence of P Cygni-like line profiles indicates moderate wind strength, and we found a best-fit mass loss rate at fixed terminal wind velocity of km s-1. We note that non-spherical wind structures, expected in Be stars and not captured by 1D models, may affect the detailed line shape and velocity centroid (e.g. Puls et al., 2008).
@@
Building on the analysis of the fundamental parameters, we further estimated the CNO surface abundances in the Be star. We found that the observed optical and UV absorption lines for these elements did not match models assuming solar abundances. Specifically, the lines for carbon and oxygen appeared weaker, while those for nitrogen appeared stronger. To better fit the observed absorption lines, we adjusted the CNO abundances in our models (see Table 3). The observed spectra are better reproduced by a model with a factor of three higher nitrogen abundance and factors of three and two lower abundances for carbon and oxygen, respectively. The CNO-processed material on the surface of the Be star could be the result of its rapid rotation, or past mass transfer from the companion star (e.g. Maeder and Meynet, 2000; Langer et al., 2008, though see also Hunter et al. 2008).